OpenStudy (anonymous):

Is this series absolutely convergent, conditionally convergent, or divergent? Sum from n=1 to infinity of (-1) ^ n * [ (3^n)/(2^n + 3^n) ] ?

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} (-1)^{n} \frac{ 3^{n} }{ 2^{n} + 3^{n} }\] I tried using the ratio test, but it does not seem to work?

OpenStudy (anonymous):

jeez this is tricky it looks like there might be two different limits depending on whether n is even or odd so it seem unlikely that it converges

OpenStudy (anonymous):

i can't get a good idea of what the partial sums are maybe there is some other method that shows it diverges

OpenStudy (anonymous):

Will the limit comparison work? I tried to compare it to \[\sum_{n=1}^{\infty} \frac{ 3^{n} }{ 3^{n} } = \sum_{n=1}^{\infty} 1\] which diverges, but then finding the limit of a_n/b_n is tricky.

OpenStudy (anonymous):

no i don't think so i am stumped

OpenStudy (anonymous):

i think the limit of the terms (not the sum) is not zero, but is one

OpenStudy (anonymous):

How are you able to determine that?

OpenStudy (anonymous):

i posted it as a question